OpenStudy (dls):

limit x->1 arcsec(2-x)/x^2

OpenStudy (dls):

can i write it as? \[\Huge \frac{1}{\cos^{-1}(2-x) \times x^2}\]

OpenStudy (dls):

actually im solving a limit,so ill use LH,now it will be 0 on seperate derivating right?

OpenStudy (dls):

@yrelhan4

OpenStudy (dls):

@phi

OpenStudy (dls):

@RadEn

OpenStudy (dls):

@.Sam.

OpenStudy (phi):

no, 1/acos(A) ≠ asec(A) in this problem you need to find A such that sec(A)=1 or 1/cos(A) = 1 then do the limit A/x^2

OpenStudy (dls):

why is it so? and this gives 0 too,and answer is 0 too. Maybe a co-incidence

OpenStudy (phi):

isn't 1/acos(1) = 1/0 = not defined ?

OpenStudy (dls):

hmm..yeah